# Inversion of adjunction for quotient singularities III: semi-invariant case

### 10 Dec 2023 v1

@@ 4. Order calculations
$\left(\bigwedge^{c}\left(I_{2}/I_{2}^{2}\right)\otimes_{S_{2}}S_{5}\right)\otimes_{S_{5}}\Omega_{S_{5}/k[t^{1/d}]}^{N-c}\to\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{5}.$
@@
The $S_{5}$-module $\left(\bigwedge^{c}\left(I_{2}/I_{2}^{2}\right)\otimes_{S_{2}}S_{5}\right)\otimes_{S_{5}}\Omega_{S_{5}/k[t^{1/d}]}^{N-c}$ is generated by the elements of the form
@@
$t^{w_{\gamma}({\bf f})}\mathfrak{n}^{\prime}_{p}\left(\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{5}\right),$
which completes the proof.
@@ 5. Minimal log discrepancies and arc spaces for quotient varieties
We prove the assertion (2) for $J=\mathfrak{n}^{\prime}_{p}$. We set
@@
For each $i=1,2,3$, let $W_{i}$ be the subscheme of $\overline{B}^{(\gamma)}$ defined by $J_{i}$. Note that $\mathfrak{n}_{r,B}\not=0$. Then, by what we have already proved, each $(W_{i})_{\infty}$ is a thin set of $\overline{B}^{(\gamma)}_{\infty}$. By Lemmas 2.12, 4.6 and 4.9, we have
@@
which implies that $W_{\infty}$ is also a thin set of $\overline{B}^{(\gamma)}_{\infty}$.
@@
Let $\mathfrak{a}\subset\mathcal{O}_{B}$ be a non-zero ideal sheaf and $\delta$ a positive real number. Then
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\delta})$ $=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\delta b_{1}\right\}$
$=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C^{\prime}_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\delta b_{1}\right\}$
@@
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\delta})=\inf_{b_{1},b_{6}\in\mathbb{Z}_{\geq 0}}\left\{\begin{array}[]{c}\operatorname{codim}\left(\operatorname{Cont}^{\geq 1}(\mathfrak{m}_{x})\cap\operatorname{Cont}^{b_{1}}(\mathfrak{a})\cap\operatorname{Cont}^{b_{6}}(\mathfrak{n}_{r,B})\right)\\ -\frac{b_{6}}{r}-\delta b_{1}\end{array}\right\}.$
@@
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\delta})$
$=\inf_{b_{1},b_{6}}\left\{\operatorname{codim}\left(\operatorname{Cont}^{\geq 1}(\mathfrak{m}_{x})\cap\operatorname{Cont}^{b_{1}}(\mathfrak{a})\cap\operatorname{Cont}^{b_{6}}(\mathfrak{n}_{r,B})\right)-\frac{b_{6}}{r}-\delta b_{1}\right\}$
$=\inf_{\gamma,{\bf b}}\left\{\operatorname{codim}\bigl{(}\overline{\mu}_{\gamma\infty}(D_{\gamma,{\bf b}})\bigr{)}-\frac{b_{6}}{r}-\delta b_{1}\right\}$
$=\inf_{\gamma,{\bf b}}\bigl{\{}\operatorname{codim}(D_{\gamma,{\bf b}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\delta b_{1}\bigr{\}}$
$=\inf_{\gamma,b_{1},b_{2}}\bigl{\{}\operatorname{codim}(C_{\gamma,b_{1},b_{2}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\delta b_{1}\bigr{\}},$
@@
The formulas in Theorem 5.3 can be easily extended to $\mathbb{R}$-ideals. For an $\mathbb{R}$-ideal sheaf $\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{\delta_{i}}$ on $B$, we have
@@
$=\inf_{\gamma\in G,w_{1},\ldots,w_{s},b\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C_{\gamma,w_{1},\ldots,w_{s},b}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b-\sum_{i=1}^{s}\delta_{i}w_{i}\right\}$
$=\inf_{\gamma\in G,w_{1},\ldots,w_{s},b\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C^{\prime}_{\gamma,w_{1},\ldots,w_{s},b}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b-\sum_{i=1}^{s}\delta_{i}w_{i}\right\}$
@@ 6. PIA formula for hyperquotient singularities
Note that $\overline{B}$ is normal by Lemma 3.2. For any closed point $x\in\overline{B}$, since $\mathcal{O}_{\overline{B},x}$ is a domain, $\mathcal{O}_{\overline{H}_{i},x}$ is also a domain by Lemma 6.2 below. This implies that $(f_{i})\cap\mathcal{O}_{A,y}$ is a prime ideal for any closed point $y\in B$, which completes the proof of (1).
Let $I_{H_{i}}:=(f_{i})\cap R^{G}\subset R^{G}$ denote the defining ideal of $H_{i}$, and let $U$ be an open neighborhood of $B$ in $A$ such that ${H_{i}}|_{U}$ is a prime Weil divisor. Then, we have $I_{H_{i}}^{d}\subset\left(f_{i}^{d}\right)\subset I_{H_{i}}$. Therefore, it is sufficient to prove $\operatorname{ord}_{{H_{i}}|_{U}}\left(f_{i}^{d}\right)=d$. Since $\overline{A}\to A$ is étale in codimension one, we have $(f_{i})R_{(f_{i})}=I_{H_{i}}R_{(f_{i})}$. Therefore, we have
$\bigl{(}f^{d}_{i}\bigr{)}R^{G}_{I_{H_{i}}}=\bigl{(}f_{i}\bigr{)}^{d}R_{(f_{i})}\cap R^{G}_{I_{H_{i}}}=I_{H_{i}}^{d}R_{(f_{i})}\cap R^{G}_{I_{H_{i}}}=I_{H_{i}}^{d}R^{G}_{I_{H_{i}}},$
@@
Let $\mathfrak{a}\subset\mathcal{O}_{A}$ be a non-zero ideal sheaf on $A$. Let $\delta$ be a non-negative real number. Let $x\in A$ be the origin. Then, we have
$\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\delta}\right)$
$=\inf_{\gamma\in G,b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}(C_{\gamma,b_{1},b_{2}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{1}-\delta b_{2}\right\},$
@@
(i) $\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\delta}\right)$
$=\inf_{\gamma\in G,b_{0},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(D_{\gamma,b_{0},b_{2}}\right)+\operatorname{age}(\gamma)-\frac{b_{0}}{d}-\delta b_{2}\right\},$
@@
Let $\mathfrak{a}\subset\mathcal{O}_{A}$ be a non-zero ideal, and let $\delta$ be a non-negative real number. Suppose that $\mathfrak{a}\mathcal{O}_{B}\not=0$. Let $x\in A$ be the origin. Suppose that $B$ is klt. Then, we have
$\operatorname{mld}_{x}\left(A,H_{1}+\cdots+H_{c},\mathfrak{a}^{\delta}\right)=\operatorname{mld}_{x}\left(B,\mathfrak{a}^{\delta}\mathcal{O}_{B}\right).$
@@
$\operatorname{mld}_{x}\left(A,H_{1}+\cdots+H_{c},\mathfrak{a}^{\delta}\right)\leq\operatorname{mld}_{x}\left(B,\mathfrak{a}^{\delta}\mathcal{O}_{B}\right).$
@@
$\operatorname{mld}_{x}\left(A,H_{1}+\cdots+H_{c},\mathfrak{a}^{\delta}\right)=\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\delta}\right).$
@@
(ii) $\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\delta}\right)$
$=\inf_{\gamma\in G,b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}_{\overline{A}^{(\gamma)}_{\infty}}(C_{\gamma,b_{1},b_{2}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{1}-\delta b_{2}\right\},$
@@
(iii) $\operatorname{mld}_{x}\left(B,\mathfrak{a}^{\delta}\mathcal{O}_{B}\right)$
$=\inf_{\gamma\in G,b_{2},b_{3}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}_{\overline{B}^{(\gamma)}_{\infty}}\left(D_{\gamma,b_{2},b_{3}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{3}-\delta b_{2}\right\},$

### 30 Apr 2026 v2

@@ 4. Order calculations
(i) $\left(\left(\bigwedge^{c}\left(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\right)\otimes_{S_{\overline{B}_{k[t^{1/d}]}}}S_{\widetilde{B}^{(\gamma)}_{1/d}}\right)\otimes_{S_{\widetilde{B}^{(\gamma)}_{1/d}}}\Omega_{S_{\widetilde{B}^{(\gamma)}_{1/d}}/k[t^{1/d}]}^{N-c}\right)$ $\otimes_{S_{\widetilde{B}^{(\gamma)}_{1/d}}}S_{\overline{B}^{(\gamma)}_{1/d}}$
$\longrightarrow\ \ \Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{\overline{B}^{(\gamma)}_{1/d}}$.
The $S_{\widetilde{B}^{(\gamma)}_{1/d}}-$module
@@
$\left(\bigwedge^{c}\left(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\right)\otimes_{S_{\overline{B}_{k[t^{1/d}]}}}S_{\widetilde{B}^{(\gamma)}_{1/d}}\right)\otimes_{S_{\widetilde{B}^{(\gamma)}_{1/d}}}\Omega_{S_{\widetilde{B}^{(\gamma)}_{1/d}}/k[t^{1/d}]}^{N-c}$
is generated by the elements of the form
@@
$t^{w_{\gamma}({\bf f})}\mathfrak{n}^{\prime}_{p}\left(\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{\overline{B}^{(\gamma)}_{1/d}}\right)$,
which completes the proof. ∎
@@ 5. Minimal log discrepancies and arc spaces for quotient varieties
We prove the assertion (2) for the case $J=\mathfrak{n}^{\prime}_{p}$. We set
@@
For each $i=1,2,3$, let $W_{i}$ be the subscheme of $\overline{B}^{(\gamma)}$ defined by $J_{i}$. Note that $\mathfrak{n}_{r,B}\neq 0$. Then, by what we have already proved, each $(W_{i})_{\infty}$ is a thin set of $\overline{B}^{(\gamma)}_{\infty}$. We shall prove the inclusion
@@
which implies that $W_{\infty}$ is also a thin set of $\overline{B}^{(\gamma)}_{\infty}$. Suppose that an arc $\alpha\in\overline{B}^{(\gamma)}_{\infty}$ satisfies $\alpha\notin(W_{1})_{\infty}\cup(W_{2})_{\infty}\cup(W_{3})_{\infty}$, that is,
  1. (ii)
    $\operatorname{ord}_{\alpha}\left(\operatorname{Jac}_{\overline{B}^{(\gamma)}/k[t]}\right)<\infty$,
  2. (iii)
    $\operatorname{ord}_{\alpha}\left(\operatorname{Jac}_{\widetilde{B}^{(\gamma)}/k[t]}\mathcal{O}_{\overline{B}^{(\gamma)}}\right)<\infty$, and
  3. (iv)
    $\operatorname{ord}_{\alpha}\left(\mathfrak{n}_{r,B}\mathcal{O}_{\overline{B}^{(\gamma)}}\right)<\infty$.
By (ii), (iv), and Lemma 4.6, we have
  1. (v)
    $\operatorname{ord}_{\alpha}\bigl(\operatorname{jac}_{\overline{\mu}_{\gamma}}\bigr)<\infty$.
Then, by (ii), (iii), (v), and Lemma 2.14 (applied to the composite morphism $\overline{\mu}_{\gamma}=\mu_{\gamma}\circ p$), we obtain
  1. (vi)
    $\operatorname{ord}_{\alpha}\left(\operatorname{jac}_{p}\right)<\infty$.
Finally, by (ii), (vi), and Lemma 4.9, we have
$\operatorname{ord}_{\alpha}(\mathfrak{n}^{\prime}_{p})<\infty,$
which implies $\alpha\notin W_{\infty}$ and finishes the proof of (i).
@@
Let $\mathfrak{a}\subset\mathcal{O}_{B}$ be a non-zero ideal sheaf and $\tau$ a positive real number. Then
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\tau})$ $=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\tau b_{1}\right\}$
$=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C^{\prime}_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\tau b_{1}\right\}$
@@
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\tau})=\inf_{b_{1},b_{6}\in\mathbb{Z}_{\geq 0}}\left\{\hskip-2.84526pt\begin{array}[]{c}\operatorname{codim}\left(\operatorname{Cont}^{\geq 1}(\mathfrak{m}_{x})\cap\operatorname{Cont}^{b_{1}}(\mathfrak{a})\cap\operatorname{Cont}^{b_{6}}(\mathfrak{n}_{r,B})\right)\\ -\frac{b_{6}}{r}-\tau b_{1}\end{array}\right\}.$
@@
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\tau})$
$=\inf_{b_{1},b_{6}}\left\{\operatorname{codim}\left(\operatorname{Cont}^{\geq 1}(\mathfrak{m}_{x})\cap\operatorname{Cont}^{b_{1}}(\mathfrak{a})\cap\operatorname{Cont}^{b_{6}}(\mathfrak{n}_{r,B})\right)-\frac{b_{6}}{r}-\tau b_{1}\right\}$
$=\inf_{\gamma,{\bf b}}\left\{\operatorname{codim}\bigl(\overline{\mu}_{\gamma\infty}(D_{\gamma,{\bf b}})\bigr)-\frac{b_{6}}{r}-\tau b_{1}\right\}$
$=\inf_{\gamma,{\bf b}}\bigl\{\operatorname{codim}(D_{\gamma,{\bf b}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\tau b_{1}\bigr\}$
$=\inf_{\gamma,b_{1},b_{2}}\bigl\{\operatorname{codim}(C_{\gamma,b_{1},b_{2}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\tau b_{1}\bigr\},$
@@
The formulas in Theorem 5.3 can be easily extended to $\mathbb{R}$-ideals. For an $\mathbb{R}$-ideal sheaf $\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{\tau_{i}}$ on $B$, we have
@@
$=\inf_{\gamma\in G,w_{1},\ldots,w_{s},b\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C_{\gamma,w_{1},\ldots,w_{s},b}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b-\sum_{i=1}^{s}\tau_{i}w_{i}\right\}$
$=\inf_{\gamma\in G,w_{1},\ldots,w_{s},b\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C^{\prime}_{\gamma,w_{1},\ldots,w_{s},b}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b-\sum_{i=1}^{s}\tau_{i}w_{i}\right\}$
@@ 6. PIA formula for hyperquotient singularities
Note that $\overline{B}$ is normal by Lemma 3.2. For any closed point $x\in\overline{B}$, since $\mathcal{O}_{\overline{B},x}$ is a domain, $\mathcal{O}_{\overline{H}_{i},x}$ is also a domain by Lemma 6.2 below. This implies that $(f_{i})\mathcal{O}_{\overline{A},x}\cap\mathcal{O}_{A,y}$ is a prime ideal for the closed point $y=q(x)\in B$, which completes the proof of (1).
Let $x\in\overline{B}$ be a closed point, and let $y:=q(x)\in B$. We define an ideal $I_{H_{i}}\subset\mathcal{O}_{A,y}$ by
$I_{H_{i}}:=(f_{i})\mathcal{O}_{\overline{A},x}\cap\mathcal{O}_{A,y}.$
By (1), $I_{H_{i}}$ is a prime ideal. Since $f_{i}$ is $G$-semi-invariant, we have $I_{H_{i}}^{d}\subset\left(f_{i}^{d}\right)\subset I_{H_{i}}$. In order to prove $\operatorname{div}\left(f_{i}^{d}\right)=dH_{i}$ on some open neighborhood of $y$, it is sufficient to show
$\bigl(f^{d}_{i}\bigr)(\mathcal{O}_{A,y})_{I_{H_{i}}}=I_{H_{i}}^{d}(\mathcal{O}_{A,y})_{I_{H_{i}}},$
where $(\mathcal{O}_{A,y})_{I_{H_{i}}}$ denotes the localization of $\mathcal{O}_{A,y}$ at the prime ideal $I_{H_{i}}$. Since $\overline{A}\to A$ is étale in codimension one, the extension of discrete valuation rings $(\mathcal{O}_{A,y})_{I_{H_{i}}}\hookrightarrow(\mathcal{O}_{\overline{A},x})_{(f_{i})}$ is unramified, which implies
$(f_{i})(\mathcal{O}_{\overline{A},x})_{(f_{i})}=I_{H_{i}}(\mathcal{O}_{\overline{A},x})_{(f_{i})}.$
Here, $(\mathcal{O}_{\overline{A},x})_{(f_{i})}$ denotes the localization of $\mathcal{O}_{\overline{A},x}$ at the prime ideal generated by $f_{i}$. Therefore, we have
$\bigl(f^{d}_{i}\bigr)(\mathcal{O}_{A,y})_{I_{H_{i}}}$ $=\bigl(f_{i}\bigr)^{d}(\mathcal{O}_{\overline{A},x})_{(f_{i})}\cap(\mathcal{O}_{A,y})_{I_{H_{i}}}$
$=I_{H_{i}}^{d}(\mathcal{O}_{\overline{A},x})_{(f_{i})}\cap(\mathcal{O}_{A,y})_{I_{H_{i}}}$
$=I_{H_{i}}^{d}(\mathcal{O}_{A,y})_{I_{H_{i}}},$
@@
Let $\mathfrak{a}\subset\mathcal{O}_{A}$ be a non-zero ideal sheaf on $A$. Let $\tau$ be a non-negative real number. Let $x\in A$ be the origin. Then, we have
$\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\tau}\right)$
$=\inf_{\gamma\in G,b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}(C_{\gamma,b_{1},b_{2}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{1}-\tau b_{2}\right\},$
@@
(i) $\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\tau}\right)$
$=\inf_{\gamma\in G,b_{0},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(D_{\gamma,b_{0},b_{2}}\right)+\operatorname{age}(\gamma)-\frac{b_{0}}{d}-\tau b_{2}\right\},$
@@
Let $\mathfrak{a}\subset\mathcal{O}_{A}$ be a non-zero ideal, and let $\tau$ be a non-negative real number. Suppose that $\mathfrak{a}\mathcal{O}_{B}\not=0$. Let $x\in A$ be the origin. Suppose that $B$ is klt. Then, we have
$\operatorname{mld}_{x}\left(A,H_{1}+\cdots+H_{c},\mathfrak{a}^{\tau}\right)=\operatorname{mld}_{x}\left(B,\mathfrak{a}^{\tau}\mathcal{O}_{B}\right).$
@@
$\operatorname{mld}_{x}\left(A,H_{1}+\cdots+H_{c},\mathfrak{a}^{\tau}\right)\leq\operatorname{mld}_{x}\left(B,\mathfrak{a}^{\tau}\mathcal{O}_{B}\right).$
@@
$\operatorname{mld}_{x}\left(A,H_{1}+\cdots+H_{c},\mathfrak{a}^{\tau}\right)=\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\tau}\right).$
@@
(ii) $\operatorname{mld}_{x}\left(A,(f_{1}^{d}\cdots f_{c}^{d})^{\frac{1}{d}}\mathfrak{a}^{\tau}\right)$
$=\inf_{\gamma\in G,b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}_{\overline{A}^{(\gamma)}_{\infty}}(C_{\gamma,b_{1},b_{2}})+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{1}-\tau b_{2}\right\},$
@@
(iii) $\operatorname{mld}_{x}\left(B,\mathfrak{a}^{\tau}\mathcal{O}_{B}\right)$
$=\inf_{\gamma\in G,b_{2},b_{3}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}_{\overline{B}^{(\gamma)}_{\infty}}\left(D_{\gamma,b_{2},b_{3}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{3}-\tau b_{2}\right\},$